OpenStudy (anonymous):

show that the equation 4x^2+4y^2-4x-12y+1=0 represents a circle. find its center and radius

OpenStudy (anonymous):

You need to complete to square for your x's and your y's to put it into standard form for a circle.

OpenStudy (slaaibak):

get in in the form: (x-a)^2 + (y-b)^2 = r^2 where r is the radius, and (a,b) the center

OpenStudy (mertsj):

First, divide each term by 4

OpenStudy (mertsj):

\[x^2-x+y^2-3y=-\frac{1}{4}\]

OpenStudy (mertsj):

Now add 1/4 and 9/4 to the x an y parts respectively and to the right side

OpenStudy (anonymous):

k i divided each term by 4 and i got \[x^2+y^2+x-3y+1/4=0\] am i ryt?

OpenStudy (mertsj):

\[x^2+x+\frac{1}{4}+y^2-3y+\frac{9}{4}=-\frac{1}{4}+\frac{1}{4}+\frac{9}{4}\]

OpenStudy (mertsj):

Now factor the two trinomial squares and add the like terms to determine the center and radius.

OpenStudy (anonymous):

k could u pls explain how u got 9/4?

OpenStudy (mertsj):

The y coefficient is 3. (1/2)(3)=3/2 and (3/2)^2=9/4

OpenStudy (anonymous):

k what do i do next to get center nd radius?

OpenStudy (mertsj):

Now factor the two trinomial squares and add the like terms to determine the center and radius.

OpenStudy (anonymous):

k i got x^2+y^2+x-3y+5/2=-5/2 am i ryt?

OpenStudy (anonymous):

mertsj how do i find center and radius from here

OpenStudy (mertsj):

\[(x+\frac{1}{2})^2+(y-\frac{3}{2})^2=\frac{9}{4}\]

OpenStudy (mertsj):

Center is (-1/2, 3/2) Radius is 3/2

OpenStudy (anonymous):

how did get it